Rocco is m years old. He his 3 years older than his brother Tony and 7 years younger than his sister Juliana. If their grandmothers age is both 5 times more than Tony's and 3 times more than Juliana's, what is Rocco's age.
Answer by ewatrrr(24785)

m is Rocco's age. Tony (m-3) and Julianna (m+7)
5(Tony) = 3(Julianna)
5(m-3) = 3(m+7)
5m - 15 = 3m + 21
2m = 36
m = 18, Rocco's age